

********************************************************************
UARS WEEKLY STATUS REPORT - 25 DEC 1998
********************************************************************
***********
SOLSTICE
**********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible. In addition, SOLSTICE, MPG, and FOT personnel prepared
to place SOLSTICE under PMON control; allowing SOLSTICE to operate for
only a short period during each spacecraft day. These operations begin
on DOY 358.
Hoping everyone had a very Happy SOLSTICE - and Happy Holidays too!
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Merry Christmas to all from the PEM team!!
PEM is waiting to conduct tests of the OBC activation. PEM activation
occurs during daylight only, when there is adequate power generated from
the UARS solar array. Currently, the PEM magnetometer and medium energy
spectrometers (particles below 30 keV) activate correctly under OBC control.
Tests conducted on the high energy particle spectrometers (HEPS -- particles
greater than 30 keV) have had limited success. New OBC activation procedure
for HEPS based on previous tests has been created. Tests of the new activation
sequence will begin shortly. Due to testing, HEPS science data from
December 1 to December 15 and after December 17 should be considered invalid
and not used.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 December 1998 (Orbit 39610, GMT Day 345) through 18
December 1998 (Orbit 39731, GMT Day 352).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 346/2316 39642 Off
PEM 350/1828 39699 Calibration
HALOE 350/2322 39702 On
PEM 352/0102 39718 Axis On
The beta angle ranged between 9.0 degrees to 16.7 degrees for this report
period. The beta angle is now increasing toward a maximum angle of 34.0
degrees on 27 December 1998 (DOY 361).
A Reverse-to-Forward Yaw Around maneuver (number 74) was successfully
executed in Orbit 39653 beginning at 347/16:01:21 and completing at
347/16:38:31 on 13 December 1998. The SSPP Instruments (ACRIM, SOLSTICE,
SUSIM) were enabled as well as HRDI and PEM in day only operations. Due to
the SA position at 269 degrees, the effective night length of the
spacecraft after the maneuver was approximately 55 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night load
bus voltage was 25.5 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 22.0 and 21.0 percent respectively.
The next (forward-to-reverse) YAW is scheduled for 11 January 1998.
Battery 3 maximum differential voltages have stayed around 180 mV since the
yaw around maneuver. The FOT consensus is that this behavior is not
dangerous: end of night load bus voltage remains acceptable, discharge
sharing remains very nearly equal, and differential voltage excursions
remain limited to the end of the power night.
A +0.5 Degree Roll Offset Maneuver (cold side down) was performed on 17
December (DOY 351) in Orbit 39713 at 351/14:54:57. The maneuver was
requested by the WINDII instrument to facilitate the observation of Polar
Mesospheric Clouds. UARS will return to level flight on 11 January 1999
(DOY 011). The exact time for the return maneuver has not been finalized.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 11 December 1998 and 18 December 1998 are provided in
Appendix D.
The clock error ranged from 0.4 to 10.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 00
% HALOE 02
% PEM 06
% WINDII 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 04
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496936
Transponder B: 98/300 - 2287.498964
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/351 show the observatory data loss to be
21 hours, 40 minutes, 27.316 seconds (an increase of 13.056 seconds since
last report period). This is a 0.0342 percent data loss which equals a
99.9658 percent data capture for the mission.
The increase of 13.056 seconds occurred over three (3) days:
7 1.824 seconds on Day 345 due to hit in data (IP);
7 2.048 seconds on Day 346 due to hit in data (IP); and
7 9.184 seconds on Day 351 due to data hit and insufficient overlap.
The official database release memo for Database 69 was completed. The FOT
has been using database 69 for operations since 25 September 1998, but
apparently never delivered the official release memorandum to the Mission
Services Library in Building 3/14. The memo will be delivered on 21
December.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Examination of operations with the extended PMON control plan continued.
It was determined that when HALOE is turned off with the Secondary Mirror
heater on, PMON does not turn off the heater. In this case, the TMON
monitoring the mirror temperature tripped as designed; PMON control resumed
normally on the low side of the thermal cycle. Nonetheless, the FOT will
recommend a procedural or programming change so that TMON cycling of the
heater is never required during normal operations, including instrument
activation and deactivation.
Development of the procedure for the clock change to include the 01 January
1999 leap second adjustment is complete.
Planning for testing of the Orbit Adjust Maneuver on the UTTS (simulator)
in December 1998 continues.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - UARS participated in a User Planning System
(UPS) Y2K simulation on 17 December 1998. To accommodate the simulation,
the back-up UPS system was configured to use the upcoming Y2K-compatible
software release. In addition, the NCC T98 simulator was connected to the
back-up UPS and configured to replicate the 1999 to 2000 roll-over. The
various FOTs were invited to perform normal scheduling operations using the
UPS and NCC test systems. The UARS FOT exercised most of the real-time and
forecast scheduling tasks, as well as generating various scheduling
reports. The only problem noticed by the FOTs was the unavailability to
generate a schedule listing which spans the century cross-over. This
problem will be rectified by the UPS development team, but it is a non-
issue to UARS anyway, since the UARS schedule listings are done on a day by
day basis..
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - The FOT completed the last two remaining
CMS Y2K test functions this week. These test functions dealt with the
Schedule SSPP processing for the 1999-to-2000 cross-over, as well as SSPP
processing that begins and ends solely in year 2000. This last bit of
testing yielded no identifiable Y2K problems with the CMS software. The
next goal on the Y2K test outline will be end-to-end testing which should
be coordinated sometime next month.
Flight Dynamics Facility (FDF) - No change.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
No change. A new release is expected in the near future.
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 11 January
1999.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 December 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39642 346/2316 98-493 WINDII Power OFF
39683 349/1707 98-494 PEM PEM PDBs
39697 350/1456 98-495 PEM PMON Act. 28,29 Disable
39698 350/1645 98-496 PEM PMON Grp 10 Disable
39699 350/1828 98-496 PEM HEPS ON
39701 350/2145 98-497 PEM PEM calibration stop
39702 350/2322 98-498 HALOE Power ON (Part-1)
39705 351/0434 98-499 HALOE Power ON (Part-2)
39718 352/0102 98-500 PEM Power ON AXIS
39728 352/1711 98-501 WINDII MUB DISA
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 December 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
NBRECBI R-HI NBTR B Current 39729
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 December 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-No clock adjust for this report period
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 December 1998
11 Dec 1998 - GMT Day 345 Beta = 9.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +78.4/+56.0 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +44.8/-5.6 +5.47/+4.19 +31.6/-14.4 25.4 1.020 21.5
3 +134.4/0.0 +4.19/+2.92 +30.4/-14.4 25.4 1.020 21.0
18 Dec 1998 - GMT Day 352 Beta = 16.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +67.2/+44.8 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-44.8 +6.11/+4.83 +32.0/-14.4 25.3 1.024 21.9
3 +190.4/+22.4 +5.15/+3.87 +31.2/-15.2 25.1 1.022 21.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 December 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-122 FOT 39639 346/1814 AP sent partial load
98-117 FOT 39643 347/0054 14170 Generic Late Acquisition
98-117 FOT 39706 351/0617 14170 Generic Late Acquisition
98-117 FOT 39709 351/1032 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-117 FOT DEC 98 See List 14170 GENERIC LATE ACQ
98-122 FOT 39639 346/18:14 CMD BUFFER PROB
SEVNTFW51